# CMoE: Fast Carving of Mixture-of-Experts for Efficient LLM Inference

### v1

@@ Abstract
Large language models (LLMs) achieve impressive performance by scaling model parameters, but this comes with significant inference overhead. Feed-forward networks (FFNs), which dominate LLM parameters, exhibit high activation sparsity in hidden neurons. To exploit this, researchers have proposed using a mixture-of-experts (MoE) architecture, where only a subset of parameters is activated. However, existing approaches often require extensive training data and resources, limiting their practicality. We propose CMoE (Carved MoE), a novel framework to efficiently carve MoE models from dense models. CMoE achieves remarkable performance through efficient expert grouping and lightweight adaptation. First, neurons are grouped into shared and routed experts based on activation rates. Next, we construct a routing mechanism without training from scratch, incorporating a differentiable routing process and load balancing. Using modest data, CMoE produces a well-designed, usable MoE from a 7B dense model within five minutes. With lightweight fine-tuning, it achieves high-performance recovery in under an hour. We make our code publicly available at https://github.com/JarvisPei/CMoE.
I Introduction
Large language models (LLMs) have demonstrated exceptional proficiency in managing complex tasks and exhibiting emergent capabilities across diverse domains and applications, particularly when scaled to billions of parameters [1, 2, 3, 4]. While LLMs have attained remarkable success, their expanding computational demands and model sizes have intensified challenges related to practical deployment, especially in environments with limited hardware resources or stringent latency requirements. To mitigate these challenges, mixture-of-experts (MoE) architectures [5, 6, 7, 8] have emerged as a promising paradigm. Unlike dense LLMs, where all parameters are activated for every input token, MoE models replace monolithic feed-forward networks (FFNs) with sparsely activated experts: specialized sub-networks that process inputs conditionally via dynamic routing. This design decouples model capacity from computational cost—activating only a subset of experts per token while preserving the model’s expressive power.
Recently, researchers have found that there is high activation sparsity in the hidden neurons of FFNs of dense LLMs, which motivates them to develop sparsity-aware acceleration techniques to reduce computational overhead while maintaining model performance [9, 10]. Building upon this insight, a growing body of research has focused on transforming dense LLMs into MoE architectures through strategic reorganization of FFN parameters but not training MoE from scratch [11, 12, 13]. The prevailing methodology replaces conventional FFN layers with MoE layers, where neurons are partitioned into multiple expert sub-networks while maintaining the original parameter count. During inference, a routing mechanism selectively activates only a subset of experts per input token, thereby achieving dynamic sparsity without compromising model capacity. However, due to the high sparsity target always required in MoE models, these works often need massive computing resources and billions of training data for continual pre-training on the constructed MoE models.
To address the aforementioned limitations, we propose carved MoE, named CMoE, a framework that efficiently carves sparse MoE architectures from dense LLMs through parameter reorganization and training-free structural adaptation. Unlike prior approaches that rebuild MoE models from scratch via resource-intensive pre-training, CMoE strategically “carves” experts from the dense model’s existing feed-forward network (FFN) neurons while preserving their inherent knowledge. This carving process is both computationally efficient—requiring only minutes of lightweight processing—and sophisticated, as it leverages systematic neuron grouping and analytical router construction to retain performance with minimal fine-tuning. The core innovation lies in CMoE’s ability to restructure dense FFNs into MoE layers without re-training the base model. First, we identify neurons that universally encode common knowledge (a.k.a. shared experts) and those that exhibit specialized, input-dependent activation patterns (a.k.a. routed experts). Shared experts are systematically retained by selecting neurons with the highest activation rates, ensuring they capture broadly applicable features. By formulating routed expert grouping as a balanced linear assignment problem, solved via the Jonker-Volgenant algorithm, CMoE clusters neurons into experts while maintaining parameter balance and activation coherence. Second, we derive the routing mechanism directly from the dense model’s activation statistics, bypassing the need for end-to-end router training. This involves constructing a differentiable routing function initialized using representative neurons from each expert cluster, enabling immediate usability while preserving optimization potential.
In summary, the key contributions of this paper are:
  • CMoE: A framework that efficiently carves MoE from dense LLMs by reorganizing FFN neurons into shared/routed experts, eliminating costly pre-training.
  • Shared and routed experts: ensuring parameter balance and activation coherence for CMoE.
  • Training-Free Routing: using representative neurons and lightweight adaptation for router initialization, which enables rapid performance recovery.
Extensive experiments show that, with activation ratio $25\%$, CMoE can maintain a reasonable perplexity even without fine-tuning, and achieve $76.59\%$ accuracies of dense models on some downstream benchmarks with lightweight fine-tuning on 2,048 samples.
II Related Work
@@
In contrast to pretraining MoE models from scratch, recent research has investigated the feasibility of constructing MoE architectures by repurposing existing dense LLMs. Current methodologies for deriving MoE models from dense checkpoints generally follow two paradigms: (1) partitioning parameters of FFNs while preserving the original model’s total parameter count [14, 10], or (2) expanding the model’s overall capacity while retaining activation dimensions comparable to standard dense models [15, 16]. This work prioritizes the former approach. Notably, MoEBERT [14] introduces an importance-driven strategy to transform FFNs into expert modules by strategically redistributing top-scoring neurons across specialized components. Concurrently, MoEfication [10] leverages the discovery of sparse activation patterns in ReLU-based FFNs within T5 architectures, enabling the decomposition of these layers into distinct expert groups governed by a learned routing mechanism. Based on continual training, the LLaMA-2 7B model is modified as a LLaMA-MoE-3.5B MoE model, where the parameters of the original FFNs are partitioned into multiple experts [11]. After training with 200B tokens, the LLaMA-MoE-3.5B model significantly outperforms dense models that contain similar activation parameters. Furthermore, based on a two-stage post-training strategy, an MoE model is constructed from the LLaMA3 8B model, where both attention and MLP are partitioned into MoE blocks [12]].
Extensive experiments have shown the effectiveness of constructing an MoE model from a dense model, and many techniques can be utilized to guarantee performance recovery. However, such performance recovery is extremely resource-consuming, which is unfavorable for efficient deployment in industrial applications. Therefore, more lightweight methods are required, such that performance recovery can be done within hours and even training-free.
@@
Note that model compression such as pruning and quantization is another important technique for efficient LLM inference [17, 18, 19, 20]. Pruning is among the most widely utilized approaches to detect and remove redundant or less significant parameters from models, thereby resulting in a sparser weight matrix and faster inference. ShortGPT [21] has put forward a simple layer-removal approach. This approach is based on block influence, which is determined by the similarity between a layer’s input and output. SliceGPT [22] substitutes each weight matrix with a smaller dense matrix, thereby decreasing the embedding dimension of the network. By taking into account contextual sparsity in real time, Deja Vu [9] and FuseGPT [23] have been proposed to accelerate LLM inference. In contrast to the post-training methods, Learn-To-be-Efficient is designed to train efficiency-aware LLMs so that they learn to activate fewer neurons and achieve a more favorable trade-off between sparsity and performance [24].
III Background
This study primarily focuses on the LLaMA family [2, 25], which uses SwiGLU [26] as the activation function. However, our analysis and findings can be adapted to most of the FFN structures of existing LLMs, including the ReLU-based FFNs [27].
@@ III Background
An FFN exists in the tail of each transformer block, which gets the input embedding $\mathbf{x}\in\mathbb{R}^{d}$ and then contributes to the output together with the residual connection, i.e. $\mathbf{x}+F(\mathbf{x})$. Typically, an FFN is a two-layer fully connected network, i.e. the up projection and down projection layer, with an activation function between them. For LLaMA, the SwiGLU composes another gate projection layer. Given the up projection weight $\mathbf{W}_{up}\in\mathbb{R}^{d\times d_{h}}$, the gate projection weight $\mathbf{W}_{gate}\in\mathbb{R}^{d\times d_{h}}$ and the down projection weight $\mathbf{W}_{down}\in\mathbb{R}^{d_{h}\times d}$, the process of an FFN is given by:
$F(\mathbf{x})=\mathbf{h}\mathbf{W}_{down},$ (1)
@@
$\mathbf{h}\!=\!\text{SwiGLU}(\mathbf{x}\mathbf{W}_{up})\!=\!\text{Swish}(\mathbf{x}\mathbf{W}_{gate})\odot(\mathbf{x}\mathbf{W}_{up}),$
@@
The basic MoE architecture is composed of a set of $N$ independent FFNs as experts, $\{E_{1},E_{2},...,E_{N}\}$, and a router network $G$ [5]. The output of an MoE-version FFN is then obtained by
$\!\!\!\!F_{MoE}(\mathbf{x})$ $=\!\sum^{N}_{i=1}g_{i}E_{i}(\textbf{x}),$ (2)
$g_{i}$ =$=\!\left\{\begin{array}[]{rl}s_{i},\!&s_{i}\in\text{TopK}(\{s_{i}|1\leq j\leq N\},K),\\ 0,\!&\text{otherwise},\end{array}\right.$
s $=\!\left[s_{1},s_{2},\cdots,s_{N}\right]=G(\textbf{x}),$
where $g_{i}$ is the score for the $i$-th expert, $\textbf{s}\in\mathbb{R}^{N}$ is the token-to-expert affinity, i.e. the output of $G$, and $\text{TopK}(\cdot,K)$ denotes the set comprising $K$ highest scores among the affinity scores calculated for $\mathbf{x}$ on all experts.
IV Methodology
[…]
Figure 1: The overview of our proposed CMoE.
IV-A Shared and Routed Experts Grouping
CMoE starts with grouping neurons in FFN into experts. Shared experts are expected to process common knowledge instead of specialization. Thus, the key idea is to group neurons that are always activated during FFN inference. For routed expert grouping, as described in previous works [28], the key idea is to cluster the neurons that are always activated simultaneously. CMoE carries forward these ideas but comes up with a detailed and analytical perspective. In this section, we construct $N$ experts of size $m=\frac{d_{h}}{N}$ (assuming $d_{h}$ is a multiple of $N$, i.e., $N\mid d_{h}$), including $N_{s}$ shared experts and $N_{r}$ routed experts $(N_{s}+N_{r}=N)$.
We begin with analyzing the isolated contribution of each neuron to the output of FFN. Consider the $i$-th element of $\mathbf{h}\in\mathbb{R}^{d_{h}}$: $h_{i}=\text{Swish}(\mathbf{x}\cdot\mathbf{w}_{gate,i})\cdot(\mathbf{x}\cdot\mathbf{w}_{up,i})$, where $\mathbf{w}_{gate,i}\in\mathbb{R}^{d}$ and $\mathbf{w}_{up,i}\in\mathbb{R}^{d}$ are the $i$-th column of $\mathbf{W}_{gate}$ and $\mathbf{W}_{up}$, respectively. It shows that the value of $h_{i}$ only depends on the input $\mathbf{x}$ and the corresponding $i$-th weights, which implies the independence between neurons. On the other hand, the output of FFN can be written as:
@@ IV-A Shared and Routed Experts Grouping
where $\mathbf{w}_{down,i}\in\mathbb{R}^{d}$ is the $i$-th row of $\mathbf{W}_{down}$. When we revisit the FFN process, we can regard each $h_{i}$ as the score to be multiplied to the split vector $\mathbf{w}_{down,i}$, whose product contributes to part of the output $F(\mathbf{x})$. As a finding of some structured pruning research [29, 30], the norm of $F(\mathbf{x})$ is always small due to the residual connection. Such phenomenon implies the high sparsity existed in FFN, and by (3) we relate it to $h_{i}$, whose value decides how large an isolated neuron contributes to the output. Therefore we make a hypothesis:
@@
which is reasonable since when $h_{i}$ is extremely small, the product $h_{i}\mathbf{w}_{down,i}$ will also vanish. It is expected that the hidden state $\mathbf{h}$ should be highly sparse, which means $|h_{i}|$ is often extremely small. To verify it, we hack into the FFN and draw the distribution of $\mathbf{h}$.
Figure 2: The histogram of FFN hidden state $\mathbf{h}$ for the $3$-th block and the $1,000$-th token.
As demonstrated in fig. 2, the distribution is sharply peaked at 00 and constrained within a small range, indicating that most $h_{i}$ are concentrated near zero and confirming the sparsity. And it exhibits symmetry, suggesting a near-normal distribution centered around the mean. Based on what we discuss and observe above, the hidden state values work well as the basis for judgment of neuron activation, because of its high differentiation and independence across different neurons. Therefore, we propose a new metric, called absolute TopK (ATopK), to determine the activation status of a neuron with index $i$:
@@
where we choose neurons with $K_{a}$ highest absolute value among the hidden state values, and assign their labels in the activation marker $\mathbf{a}=\left[a_{1},a_{2},\cdots,a_{d_{h}}\right]$ with $1$.
To further evaluate the activation information, we make samples in the training set to record their activation markers. Given a batched input tensor $\mathbf{X}\in\mathbb{R}^{b\times s\times d}$, with $b$ the batch size and $s$ the sequence length, we obtain the batched hidden state $\mathbf{H}\in\mathbb{R}^{b\times s\times d_{h}}$ as follows:
Note that in practical implementation, we normalize $\mathbf{X}$, $\mathbf{W}_{gate}$ and $\mathbf{W}_{up}$ before the calculation to eliminate the influence of their magnitudes on the output. We then calculate the activation markers for the hidden state of all the tokens, which is reshaped as an activation feature matrix $\mathbf{A}\in\mathbb{R}^{q\times d_{h}},q=b\cdot s$. Denote the $i$-th column of $\mathbf{A}$ as the feature vector $\mathbf{c}_{i}\in\mathbb{R}^{q}$, i.e., $\mathbf{A}=[\mathbf{c}_{1}\;\mathbf{c}_{2}\;\cdots\;\mathbf{c}_{d_{h}}]$ which represents the $i$-th neuron’s activation status on the sampled $q$ tokens. By calculating the expected value of each $\mathbf{c}_{i}$, we can obtain the vector of activated rates $\bm{\mu}$ as:
$\bm{\mu}=[\mu_{1},\mu_{2},\cdots,\mu_{d_{h}}],\mbox{ with }\mu_{i}=\frac{1}{q}\sum_{j=1}^{q}c_{j,i},$ (7)
where $c_{j,i}$ indicates the $j$-th element of $\mathbf{c}_{i}$. We draw the histogram of $\bm{\mu}$ as in Fig. 3.
Figure 3: The histogram of activation rates $\bm{\mu}$ for the $3$-th block with $K_{a}=1,000$.
The histogram reveals a highly skewed distribution of activation rates, where the majority of neurons exhibit low activation rates (below 0.1), with a sharp peak near 0.07. However, the distribution also features a long tail, indicating the presence of a subset of neurons with significantly higher activation rates extending up to 1. These high-activation neurons are likely active across a wide range of input tokens, making them suitable for processing common knowledge rather than task-specific specialization. Therefore, we construct shared experts by grouping these high-activation neurons. Given the total number of shared experts as $N_{s}$ and the expert size $m$, we get the selection indices set $S_{N_{s}}$ by selecting $N_{s}\cdot m$ neurons with highest activation rates:
Since the independency across neurons, as we discussed, we build the experts using the same architecture as the original FFN and assign the expert parameters to the neurons of selected indices. Since all shared experts are always activated, we just construct them together for convenience:
$\mathbf{W}_{up}^{s}$ $=\mathbf{W}_{up}[:,S_{N_{s}}],$ (9)
$\mathbf{W}_{gate}^{s}$ $=\mathbf{W}_{gate}[:,S_{N_{s}}],$
$\mathbf{W}_{down}^{s}$ $=\mathbf{W}_{down}[S_{N_{s}},:],$
where $\mathbf{W}_{up}^{s}\in\mathbb{R}^{d\times N_{s}\cdot m}$, $\mathbf{W}_{gate}^{s}\in\mathbb{R}^{d\times N_{s}\cdot m}$ and $\mathbf{W}_{down}^{s}\in\mathbb{R}^{N_{s}\cdot m\times d}$ are the weights of linear layers of the shared experts $E^{s}$. Given the input embedding $\mathbf{x}$, the output of $E^{s}$ is obtained via
@@
$E^{s}(\mathbf{x})\!=\!\mathbf{h}^{s}\mathbf{W}_{down}^{s},\mbox{with }\mathbf{h}^{s}\!=\!\text{Swish}(\mathbf{x}\mathbf{W}_{gate}^{s})\!\odot\!(\mathbf{x}\mathbf{W}_{up}^{s}).$
The majority of low activation rates also encourage us to construct routed experts, which are not always activated but are specialized for tokens encountered. To construct $N_{r}$ routed experts, we develop a customized balanced K-means clustering algorithm. We first identify $N_{r}$ centroids as the neurons (excluding those assigned to shared experts) with the highest activation rates, and group their feature vectors into a centroid set $C=\{\mathbf{c}_{i}:\mu_{i}\in\text{TopK}(\{\mu_{j}\mid 1\leq j\leq d_{h},j\notin S_{Ns}\},N_{r})\}=\{\hat{\mathbf{c}}_{1},\dots,\hat{\mathbf{c}}_{N_{r}}\}$, where, for convenience, we re-label the centroids with $\hat{\mathbf{c}}$.
By selecting these centroids, we ensure that the clusters are initialized around meaningful and prominent activation patterns, providing a strong starting point for clustering. Using the centroids selected, we will construct clusters with remaining neurons. We pay attention to their correlation to the centroids, i.e. we study the distances of their feature vectors to the centroids’. The feature vector $\mathbf{c}$ individually contains the activation status of a specific neuron, and we expect that neurons with similar functionalities will have similar patterns on the activation status, i.e. they are often activated and deactivated together during inference. Therefore, we compute the $L_{2}$ distance of each feature vector $\mathbf{c}$ to each centroid $\hat{\mathbf{c}}$. Define a distance matrix $\mathbf{D}\in\mathbb{R}^{N_{r}\cdot m\times N_{r}}$, whose element $d_{i,j}$ is the $L_{2}$ distance between the $i$-th feature vector $\mathbf{c}_{i}$ and the $j$-th centroid $\hat{\mathbf{c}}_{j}$:
$d_{i,j}=\|\mathbf{c}_{i}-\hat{\mathbf{c}}_{j}\|_{2}=\sqrt{\sum_{k=1}^{q}(c_{k,i}-\hat{c}_{k,j})^{2}},$ (10)
where, for convenience, we also include the original feature vectors of centroids inside the set to be grouped.
@@
To group the neurons into routed experts, with the centroids designed above at $t=0$, the constrained balanced $K$-means clustering algorithm works as follows [31]. Given $N_{r}\cdot m$ vectors $\mathbf{c}$, the cluster membership value $m$ and cluster centroids $\hat{\mathbf{c}}_{1}^{t},\hat{\mathbf{c}}_{2}^{t},\dots,\hat{\mathbf{c}}_{N_{r}}^{t}$ at iteration $t$, compute $\hat{\mathbf{c}}_{1}^{t+1},\hat{\mathbf{c}}_{2}^{t+1},\dots,\hat{\mathbf{c}}_{N_{r}}^{t+1}$ at iteration $t+1$ with the following $2$ steps:

### v2

# CMoE: Converting Mixture-of-Experts from Dense to Accelerate LLM Inference

@@ Abstract
Scaling large language models (LLMs) improves performance but dramatically increases inference costs. The feed-forward network (FFN), consuming approximately 70% of inference compute, represents a critical bottleneck, particularly in large batch size scenarios. While mixture-of-experts (MoE) architectures leverage activation sparsity for efficiency, converting existing dense models to MoEs traditionally requires resource-intensive continual pre-training. We present CMoE, a framework that rapidly transforms dense LLMs into MoEs without training. The key innovation lies in analyzing FFN neuron activations to partition them into shared (always active) and routed experts. Routed neurons are clustered using a balanced assignment algorithm, and a differentiable router is constructed analytically from activation statistics, enabling immediate deployment or optional lightweight fine-tuning. Experiments demonstrate that, with activation ratio of 75%, it achieves remarkable results, delivering lossless precision in terms of perplexity while still maintaining a 5% acceleration. Further experiments reveal that a CMoE configuration activating just 25% of parameters reduces end-to-end latency by 1.5× while preserving usable perplexity without additional training. Moreover, a brief LoRA fine-tuning process (requiring only 1 hour and 2,000 samples) successfully recovers over 76% of the dense model’s downstream accuracy. By effectively balancing performance and efficiency, CMoE offers a viable path forward for deploying LLMs in real-world scenarios where computational resources are limited. We make our code publicly available at https://github.com/JarvisPei/CMoE.
1 Introduction
Large language models (LLMs) achieve remarkable performance zhang2022opt ; touvron2023llama ; liu2024visual ; liu2024deepseek , but their growing size presents deployment challenges due to high computational demands, especially on resource-constrained hardware or under strict latency budgets. The mixture-of-experts (MoE) architecture lepikhin2020gshard ; du2022glam ; fedus2022switch ; dai2024deepseekmoe offers a path forward. Instead of activating all parameters per token like dense models, MoEs use sparse feed-forward networks (FFNs) composed of specialized sub-networks (experts) selected by a dynamic router. This decouples model capacity from computational cost.
In modern LLMs such as the Llama-series and DeepSeek-series, FFN blocks are responsible for approximately 70% of both inference FLOPs and memory traffic. As batch sizes increase, a common strategy for maximizing hardware utilization and throughput, models increasingly operate in a compute-bound regime, where computational power becomes the limiting factor for inference speed.
Given the FFN’s disproportionate contribution to memory and computational demands, sparsifying FFNs has become a primary target for efficiency improvement. Several studies have reported high activation sparsity in FFN neurons dejavu ; moefication ; pei2024fusegpt , inspiring a wave of techniques that aim to restructure dense FFN layers into sparse mixtures of experts (MoEs) by reorganizing existing parameters llama-moe ; llama-moe-v2 ; zheng2024learn . These MoE conversions can dramatically reduce both memory usage and computational overhead, improving throughput, particularly in large-batch, latency-sensitive inference scenarios.
However, despite their promise, such methods generally require substantial additional computation and large-scale datasets for continual pre-training or fine-tuning after conversion, in order to recover or maintain model quality. This retraining step can be prohibitively expensive and slow, posing a significant barrier to rapid deployment in production environments. As a result, there is growing interest in developing lightweight strategies that minimize retraining costs while delivering tangible inference speedups in real-world settings.
To overcome this bottleneck, we propose CMoE, a framework to efficiently convert dense LLMs into MoE architectures without requiring any training. The conversion process leverages the existing knowledge within the dense model’s FFN and completes in under five minutes on a single GPU. CMoE first analyzes neuron activation patterns across a small calibration dataset to distinguish frequently active neurons (grouped into shared experts) from sparsely active ones (grouped into routed experts). Routed experts are formed by clustering neurons using a balanced assignment algorithm jonker1988shortest to maintain activation coherence. Crucially, CMoE then constructs a differentiable routing mechanism analytically from activation statistics, bypassing the need for router training and enabling immediate use or subsequent lightweight fine-tuning.
Our contributions are:
  • Training-free conversion. CMoE rapidly restructures dense FFNs into shared and routed experts using activation analysis and balanced assignment, requiring only minutes and no gradient updates. In practical implementation, it is found that with an activation ratio of 75%, it achieves remarkable results, delivering lossless precision in terms of perplexity while still maintaining a 5% acceleration.
  • Analytical, differentiable router. The router is initialized directly from activation statistics, providing immediate functionality and a strong starting point for optional fine-tuning.
  • Controllable performance-efficiency trade-off. CMoE allows adjusting the expert activation ratio. Experiments show a spectrum from 1.5×1.5$\times$1.5 × latency reduction (at 25% activation) to near-dense perplexity (at 75% activation). Minimal fine-tuning (1 hour, 2k samples) recovers >76% of dense model downstream accuracy.
CMoE’s combination of speed, flexibility, and training-free nature presents a compelling solution for researchers and practitioners requiring efficient, low-overhead deployment of LLMs.
2 Related Work
@@
In contrast to pretraining MoE models from scratch, recent research has investigated the feasibility of constructing MoE architectures by repurposing existing dense LLMs. Current methodologies for deriving MoE models from dense checkpoints generally follow two paradigms: (1) partitioning parameters of FFNs while preserving the original model’s total parameter count zuo2022moebert ; moefication ; yang2024xmoe , or (2) expanding the model’s overall capacity while retaining activation dimensions comparable to standard dense models komatsuzaki2022sparse ; wu2024parameter . This work prioritizes the former approach. Notably, MoEBERT zuo2022moebert introduces an importance-driven strategy to transform FFNs into expert modules by strategically redistributing top-scoring neurons across specialized components. Concurrently, MoEfication moefication leverages the discovery of sparse activation patterns in ReLU-based FFNs within T5 architectures, enabling the decomposition of these layers into distinct expert groups governed by a learned routing mechanism. Based on continual training, llama-moe modifies llama-moe modifies the LLaMA-2 7B model as a LLaMA-MoE-3.5B MoE model, where the parameters of the original FFNs are partitioned into multiple experts. In llama-moe-v2 , based on a two-stage post-training strategy, an MoE model is constructed from the LLaMA3 8B model, where both attention and MLP are partitioned into MoE blocks.
@@
Note that model compression such as pruning and quantization is another important technique for efficient LLM inference lin2024awq ; pei2023quantization ; zou2024bie ; lin2024qserve ; lu2024not . Pruning is among the most widely utilized approaches to detect and remove redundant or less significant parameters from models, thereby resulting in a sparser weight matrix and faster inference. ShortGPT shortgpt determines the layer-removal by the similarity between a layer’s input and output. SliceGPT slicegpt substitutes each weight matrix with a smaller dense matrix oo deease the embedding dimension. By taking into account contextual sparsity in real time, Deja Vu dejavu has been proposed to accelerate LLM inference. In contrast to the post-training methods, Learn-To-be-Efficient is designed to train efficiency-aware LLMs so that they learn to activate fewer neurons and achieve a more favorable trade-off between sparsity and performance.
3 Methodology
Figure 1: The overview of our proposed CMoE.
CMoE transforms dense LLMs into sparsely activated MoE architectures through two key phases: efficient expert grouping and training-free router construction, followed by optional lightweight adaptation. As shown in Figure 1(a), CMoE operates through the following systematic process:
A. Neuron Activation Profiling (Section 3.1) Using a small calibration dataset, CMoE profiles the activation patterns of neurons within each FFN layer to categorize them into shared experts (high-activation, task-agnostic) and routed experts (sparsely activated, task-specific).
B. Expert Grouping (Section 3.1) Shared Experts: Neurons exhibiting the highest activation rates are directly assigned to shared experts, which remain consistently activated during inference. Routed Experts: The remaining neurons are efficiently partitioned into routed experts through balanced clustering, mathematically formulated as a linear assignment problem.
C. Router Construction (Section 3.2) The routing mechanism is analytically derived from the activation statistics of representative neurons in each expert cluster, effectively eliminating the need for computationally expensive end-to-end training.
We further enhance CMoE by making the routing function differentiable (Section 3.3), facilitating further alignment and performance recovery by optional fine-tuning when resources permit.
3.1 Shared and Routed Experts Grouping
@@
Note that an FFN exists in the tail of each transformer block, which gets the input embedding $\mathbf{x}\in\mathbb{R}^{d}$ and then contributes to the output together with the residual connection, i.e. $\mathbf{x}+F(\mathbf{x})$. Typically, an FFN is a two-layer fully connected network, i.e. the up projection and down projection layers, with an activation function between them. For LLaMA, the SwiGLU composes another gate projection layer. Given the up projection weight $\mathbf{W}_{up}\in\mathbb{R}^{d\times d_{h}}$, the gate projection weight $\mathbf{W}_{gate}\in\mathbb{R}^{d\times d_{h}}$ and the down projection weight $\mathbf{W}_{down}\in\mathbb{R}^{d_{h}\times d}$, the process of an FFN is given by:
@@
$F(\mathbf{x})=\mathbf{h}\mathbf{W}_{down},\mbox{ and }\;\mathbf{h}=\text{SwiGLU}(\mathbf{x}\mathbf{W}_{up})=\text{Swish}(\mathbf{x}\mathbf{W}_{gate})\odot(\mathbf{x}\mathbf{W}_{up}),$ (1)
@@
CMoE starts with grouping neurons in FFNs into experts. We construct $N$ experts of size $m=\frac{d_{h}}{N}$ (assuming the intermediate size $d_{h}$ is a multiple of $N$, i.e., $N\mid d_{h}$), including $N_{s}$ shared experts and $N_{r}$ routed experts $(N_{s}+N_{r}=N)$.
Shared experts are expected to process common knowledge instead of specialization. Thus, the key idea is to group neurons that are always activated during FFN inference. Given $\mathbf{x}\in\mathbb{R}^{d}$ the input embedding, consider the $i$-th element of hidden state $\mathbf{h}\in\mathbb{R}^{d_{h}}$: $h_{i}=\text{Swish}(\mathbf{x}\cdot\mathbf{w}_{gate,i})\cdot(\mathbf{x}\cdot\mathbf{w}_{up,i})$, where $\mathbf{w}_{gate,i}\in\mathbb{R}^{d}$ and $\mathbf{w}_{up,i}\in\mathbb{R}^{d}$ are the $i$-th column of $\mathbf{W}_{gate}\in\mathbb{R}^{d\times d_{h}}$ (gate projection weight) and $\mathbf{W}_{up}\in\mathbb{R}^{d\times d_{h}}$ (up projection weight), respectively. It shows that the value of $h_{i}$ only depends on the input $\mathbf{x}$ and the corresponding $i$-th weights, which implies the independence between neurons. On the other hand, the output of FFN can be written as:
@@
where $\mathbf{w}_{down,i}\in\mathbb{R}^{d}$ is the $i$-th row of $\mathbf{W}_{down}\in\mathbb{R}^{d_{h}\times d}$ (down projection weight). When we revisit the FFN process, we can regard each $h_{i}$ as the score to be multiplied to the split vector $\mathbf{w}_{down,i}$, whose product contributes to part of the output $F(\mathbf{x})$. Therefore, we make a hypothesis:
@@
which is reasonable since when $h_{i}$ is extremely small, the product $h_{i}\mathbf{w}_{down,i}$ will also vanish. In Appendix A, we conduct experiments to demonstrate that the hidden state $\mathbf{h}$ is usually highly sparse, which means $|h_{i}|$ is often extremely small. Therefore, we propose a new metric, called absolute TopK (ATopK), to determine the activation status of a neuron with index $i$:
@@
where we choose neurons with $K_{a}$ highest absolute value among the hidden state values, and assign their labels in the activation marker $\mathbf{a}=\left[a_{1},a_{2},\cdots,a_{d_{h}}\right]$ with $1$. We then obtain the set of selection indices $S_{N_{s}}$ by selecting $N_{s}\cdot m$ neurons with the highest activation rates. And we construct the shared experts as follows:
$\mathbf{W}_{up}^{s}=\mathbf{W}_{up}[:,S_{N_{s}}];\;\;\mathbf{W}_{gate}^{s}=\mathbf{W}_{gate}[:,S_{N_{s}}];\;\;\mathbf{W}_{down}^{s}=\mathbf{W}_{down}[S_{N_{s}},:];$ (5)
@@
$\Longrightarrow E^{s}(\mathbf{x})=\mathbf{h}^{s}\mathbf{W}_{down}^{s},\mbox{ with }\;\mathbf{h}^{s}=\text{Swish}(\mathbf{x}\mathbf{W}_{gate}^{s})\!\odot\!(\mathbf{x}\mathbf{W}_{up}^{s}).$
Here, $\mathbf{W}_{up}^{s}\in\mathbb{R}^{d\times N_{s}\cdot m}$, $\mathbf{W}_{gate}^{s}\in\mathbb{R}^{d\times N_{s}\cdot m}$ and $\mathbf{W}_{down}^{s}\in\mathbb{R}^{N_{s}\cdot m\times d}$ are the weights of linear layers of the shared experts $E^{s}$. More detailed analysis is given in Appendix A.
To construct $N_{r}$ routed experts, the key idea is to cluster the neurons that are always activated simultaneously. We first identify $N_{r}$ centroids by selecting neurons (excluding those already assigned to shared experts) with the highest activation rates. We group their feature vectors into a centroid set $C=\{\mathbf{c}_{i}:\mu_{i}\in\text{TopK}({\mu_{j}\mid 1\leq j\leq d_{h},j\notin S_{N_{s}}},N_{r})\}=\{\hat{\mathbf{c}}_{1},\dots,\hat{\mathbf{c}}_{N_{r}}\}$, where $\boldsymbol{\mu}=[\mu_{1},\mu_{2},\cdots,\mu_{d_{h}}]$ is the activation rate averaged from the marker $\mathbf{a}$ and we relabel the centroids as $\hat{\mathbf{c}}$. Each feature vector $\mathbf{c}$ encapsulates the activation status of a specific neuron across the calibration dataset. We hypothesize that neurons with similar functionalities exhibit similar activation patterns—specifically, they tend to be activated and deactivated together during inference. Based on this principle, we compute the $L_{2}$ distance between each feature vector $\mathbf{c}$ and each centroid $\hat{\mathbf{c}}$.
We formalize this by constructing a distance matrix $\mathbf{D}\in\mathbb{R}^{N_{r}\cdot m\times N_{r}}$, where element $d_{i,j}$ represents the $L_{2}$ distance between the $i$-th feature vector $\mathbf{c}_{i}$ and the $j$-th centroid $\hat{\mathbf{c}}_{j}$:
$d_{i,j}=\|\mathbf{c}_{i}-\hat{\mathbf{c}}_{j}\|_{2}=\sqrt{\sum_{k=1}^{q}(c_{k,i}-\hat{c}_{k,j})^{2}}.$ (6)
@@
To group the neurons into routed experts, with the centroids designed above at $t=0$, we adopt the constrained balanced $K$-means clustering method. Given $mN_{r}$ vectors $\mathbf{c}$, the cluster membership value $m$ and cluster centroids $\hat{\mathbf{c}}_{1}^{t},\hat{\mathbf{c}}_{2}^{t},\dots,\hat{\mathbf{c}}_{N_{r}}^{t}$ at iteration $t$, compute $\hat{\mathbf{c}}_{1}^{t+1},\hat{\mathbf{c}}_{2}^{t+1},\dots,\hat{\mathbf{c}}_{N_{r}}^{t+1}$ at iteration $t+1$ with the following $2$ steps: